# Fair Column Subset Selection

### v3

@@ Abstract.
Column Subset Selection, Dimensionality Reduction, Algorithmic Fairness, Matrix Factorization, Low Rank Approximation
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3672005isbn: 979-8-4007-0490-1/24/08ccs: Theory of computation Unsupervised learning and clusteringccs: Computing methodologies Feature selectionccs: Mathematics of computing Dimensionality reduction
1. Introduction
@@ 2. Related Work
Algorithmic fairness. The influential work of Dwork et al. (2012) established a formal notion of fairness in algorithmic decision-making, which has served as a foundation for subsequent research in the field. Pedreschi et al. (2008) addressed discrimination in data mining, while (Kamiran and Calders, 2010, 2012) proposed a framework for mitigating discrimination in classification. Since then, there has been extensive research on algorithmic fairness from various disciplines, including economics, game theory, statistics, ethics, and computer science (Rambachan et al., 2020; Sætra et al., 2022; Raji and Buolamwini, 2022; Buolamwini and Gebru, 2018). Fair objectives have been introduced into many classical computer science problems (Chierichetti et al., 2017; Samadi et al., 2018; Anagnostopoulos et al., 2020; Ghadiri et al., 2021; Froese et al., 2022; Zemel et al., 2013; Zafar et al., 2017; Zehlike et al., 2017). Fair clustering share similarities to fair CSS, with a key distinction: the former asks for a subset of fair representative data points (rows), while the later focuses on finding a fair representative column subset (attributes) (Chierichetti et al., 2017; Chhabra et al., 2021; Ghadiri et al., 2021). For further reading see (Pessach and Shmueli, 2022; Mitchell et al., 2021; Kleinberg et al., 2018).
@@ 3. Problem statement
$\min_{\begin{subarray}{c}C\in\mathbb{R}^{m\times k},\\ C\subset M\end{subarray}}\max\biggl{\{}\frac{\|A-C_{A}C_{A}^{+}A\|_{F}}{\|A-A_{k}\|_{F}},\frac{\|B-C_{B}C_{B}^{+}B\|_{F}}{\|B-B_{k}\|_{F}}\biggl{\}}.$
@@
4. Pairs of leverage scores
In this section, we discuss leverage scores in the context of the two-group fair setting. Leverage scores are a pivotal concept that is extensively studied in CSS literature, as they provide valuable insights for selecting column subsets with provable approximation guarantees. We begin by discussing the results of (Papailiopoulos et al., 2014) on leverage score sampling. Next, we show that sampling a subset of columns of minimum size in the two-group fair setting is NP-hard. Finally, we present an approximation algorithm that samples $1.5$ times the size of an optimal solution. An useful concept that is extensively studied in CSS literature is leverage scores, which is defined as follows:
@@ 4. Pairs of leverage scores
$\|A-C_{A}C_{A}^{+}A\|_{F}$ $\leq(1-\epsilon)^{-1/2}\|A-A_{k}\|_{F}\implies$
$Nloss_{A}(M,C)$ $\leq(1-\epsilon)^{-1/2}.$
@@ 5. Fair QR decompositions
Column pivoting involves finding a permutation matrix $\Pi$ for a given matrix $M$, such that $\|R_{22}\|_{F}$ is minimised. When this comes with certain guarantees, it leads to a rank revealing QR decomposition (RRQR), which forms the basis for various algorithms with approximation guarantees in CSS. Specifically, if we denote the first $k$ columns of the permutation matrix $\Pi$ as $\Pi_{k}$, then choosing the column subset $C=M\Pi_{k}$ guarantees that $loss(M,C)$ is bounded i.e., $\|M-P_{C}M\|_{F}=\|R_{22}\|_{F}$ (Boutsidis et al., 2009). We tailor Low-RRQR (L-RRQR), originally introduced by (Chan and Hansen, 1994), to accommodate the two-group fair setting.
@@
Fair pivoting. Note that L-RRQR may introduce unfairness, since we factorize the matrix $M$, without considering the error of the two groups, $A$ and $B$ separately. So the pivoted columns may benefit only one group. To address this, we adapt the pivoting strategy of L-RRQR to benefit the group that suffers the worse reconstruction error. Thus, we perform simultaneous RRQR on $A$ and $B$ and at step-$i$ obtain the corresponding $Q^{A}(i),R^{A}(i)$ and $Q^{B}(i),R^{B}(i)$. We inspect the spectra of both $R^{A}(i)$ and $R^{B}(i)$ and choose the pivot based on the following strategy: we choose the right singular vector $v$ of either $R^{A}(i)$ or $R^{B}(i)$ corresponding to $\max\{\sigma_{1}(R^{A}_{22}(i)),\sigma_{1}(R^{B}_{22}(i))\}$, and select $\Pi$ such that $|(\Pi^{T}v)|_{1}=\|v\|_{\infty}$. The algorithm pseudocode can be seen in Algorithm 2.
@@ 6. Experiments
Datasets. We use juvenile recidivism data (recidivism) from Catalunya (Tolan et al., 2019) and medical expenditure survey data 2015 (meps(Cohen et al., 2009), as well as various datasets from the UCI-ML repository (Dua and Graff, 2017):“heart-cleveland” (heart), “adult” (adult), “german-credit” (german), “credit-card” (credit), “student performance” (student), “compas-recidivism” (compas), “communities” (communities). Data is processed by removing protected attributes, converting categorical variables to one-hot encoding and normalizing each column to unit $L^{2}$-norm. Group membership is based on Sex, except for communities where group membership is majority white or a non-white community. Dataset statistics are reported in Table 1.
Table 1. Dataset statistics. $m_{A}$ and $m_{B}$ are the number of instances in groups $A$ and $B$, respectively. $n$ is the number of columns. $\gamma$($A$), $\gamma$($B$) is rank of $A$, $B$.
@@ 6.1. Algorithms Evaluation
Figure 1. MinMaxLoss for different values of $c$ and fixed target rank $k$
@@
Table 2. Performance comparison of algorithms.
@@
Table 2 reports the results for the seven largest datasets for various values of $k$. We set $\theta=k-\frac{1}{2}$ in all cases, except meps, where $\theta=\frac{3k}{4}$ to reduce the number of sampled columns. Thus in all datasets (except meps) the resulting number of columns in the first stage, is a $\sqrt{2}$-approximation to the optimal number of columns $c$. Column “c” in Table 2 indicates the number of columns returned in the sampling stage. Note that in some cases the number of columns is required to satisfy $\theta$ is significantly large. For each experiment, the best performing algorithm is highlighted in bold.
@@
Figure 2. Comparison of reconstruction error of CSS and FairCSS for groups $A$ and $B$.
@@
Figure 3. Price of fairness.
6.2. Evaluation of the fair CSS objective
We examine the imbalance in reconstruction errors of the “vanilla” CSS objective of Problem 1 and FairCSS-MinMax objective for groups $A$ and $B$. We compute the optimal solution for each objective for various $k$ through exhaustive enumeration. ${\sf\small opt}(M),{\sf\small opt}(A)$ and ${\sf\small opt}(B)$ denote the reconstruction error of vanilla CSS for matrices $M,A$ and $B$, respectively. ${\sf\small minmax}(M)={\sf\small minmax}(A,B)$ denote the reconstruction error of matrix $M$. Further, ${\sf\small fair}(A)$ and ${\sf\small fair}(B)$ denote the reconstruction errors of FairCSS-MinMax for groups $A$ and $B$, respectively. Lastly, ${\sf\small fair}(M)$ denote the reconstruction error of $M$ corresponding to the optimal solution of FairCSS-MinMax. We note that brute-force enumeration is expensive, even after parallelization and extensive optimization, so we only report results for heart, student and german datasets.
Fairness of solutions. In Figure 2 we compare ${\sf\small opt}(A)$, ${\sf\small opt}(B)$, that is, the optimal solution of vanilla CSS, and the optimal solution of FairCSS-MinMax ${\sf\small fair}(A),{\sf\small fair}(B)$. In most instances, we observe that the reconstruction errors of groups are disproportionate in vanilla CSS. However, the degree of imbalance is not uniform across datasets. One source of imbalance could be vastly different group sizes. However, as observed in student, $|A|>|B|$, but ${\sf\small opt}(A)<{\sf\small opt}(B)$. This further supports the need for sophisticated approaches to fairness in CSS, beyond mere normalization.
Price of fairness. Next, we verify how the fairness objective influences the quality of the solution in terms of reconstruction error. In Figure 3, we report the optimal solution of vanilla CSS and FairCSS-MinMax to assess the trade-off between fairness and reconstruction error. This analysis quantifies the extent to which we sacrifice reconstruction error to achieve our fairness objective. In most cases, we observe no significant difference in reconstruction error between vanilla CSS and FairCSS-MinMax, that is, ${\sf\small opt}(M)$ and ${\sf\small fair}(M)$, even though the value of ${\sf\small minmax}(M)$ is significantly higher than ${\sf\small opt}(M)$. Note that, the observations cannot be generalized across all datasets, and we cannot conclusively claim that there is no trade-off between fairness and reconstruction error. For instance, in the case of the heart dataset, we observe a significant difference between values of ${\sf\small opt}(M)$ and ${\sf\small fair}(M)$ for $k=5$.
7. Conclusion
@@ Acknowledgements
Suhas Thejaswi acknowledges support from the European Research Council (ERC) under the European Union’s Horizon $2020$ research and innovation program (grant agreement No. $945719$) and the European Unions’s SoBigData++ Transnational Access Scholarship.
Antonis Matakos acknowledges support from the Academy of Finland through the grant ”Model Management Systems: Machine learning meets Database Systems”- MLDB (32511).
@@ Appendix B Decay of Leverage Scores
Figure 4. Leverage scores of $A$ and $B$ for Table 2

### v4

@@ Abstract
1 Introduction
@@ 2 Related Work
Algorithmic fairness. The influential work of Dwork et al. (2012) established a formal notion of fairness in algorithmic decision-making, which has served as a foundation for subsequent research in the field. Pedreschi et al. (2008) addressed discrimination in data mining, while Kamiran and Calders (2010, 2012) proposed a framework for mitigating discrimination in classification. Since then, there has been extensive research on algorithmic fairness from various disciplines, including economics, game theory, statistics, ethics, and computer science (Rambachan et al., 2020; Sætra et al., 2022; Raji and Buolamwini, 2022; Buolamwini and Gebru, 2018). Fair objectives have been introduced into many classical computer science problems (Chierichetti et al., 2017; Samadi et al., 2018; Anagnostopoulos et al., 2020; Ghadiri et al., 2021; Froese et al., 2022; Zemel et al., 2013; Zafar et al., 2017; Zehlike et al., 2017). Fair clustering share similarities to fair CSS, with a key distinction: the former asks for a subset of fair representative data points (rows), while the later focuses on finding a fair representative column subset (attributes) Chierichetti et al. (2017); Chhabra et al. (2021); Ghadiri et al. (2021). For further reading see Pessach and Shmueli (2022); Mitchell et al. (2021); Kleinberg et al. (2018).
@@ 3 Problem statement
$\min_{\begin{subarray}{c}C\in\mathbb{R}^{m\times k}\\ C\subset M\end{subarray}}\max\biggl{\{}\frac{\|A-C_{A}C_{A}^{+}A\|_{F}}{\|A-A_{k}\|_{F}},\frac{\|B-C_{B}C_{B}^{+}B\|_{F}}{\|B-B_{k}\|_{F}}\biggl{\}}.$
@@
4 Pairs of leverage scores
In this section, we discuss leverage scores in the context of the two-group fair setting. Leverage scores are a pivotal concept that is extensively studied in CSS literature, as they provide valuable insights for selecting column subsets with provable approximation guarantees. We begin by discussing the results of Papailiopoulos et al. (2014) on leverage score sampling. Next, we show that sampling a subset of columns of minimum size in the two-group fair setting is NP-hard. Finally, we present an approximation algorithm that samples $1.5$ times the size of an optimal solution. An useful concept that is extensively studied in CSS literature is leverage scores, which is defined as follows:
@@ 4 Pairs of leverage scores
$\|A-C_{A}C_{A}^{+}A\|_{F}\leq(1-\epsilon)^{-1/2}\|A-A_{k}\|_{F}\implies Nloss_{A}(M,C)\leq(1-\epsilon)^{-1/2}.$
@@ 5 Fair QR decompositions
Column pivoting involves finding a permutation matrix $\Pi$ for a given matrix $M$, such that $\|R_{22}\|_{F}$ is minimised. When this comes with certain guarantees, it leads to a rank revealing QR decomposition (RRQR), which forms the basis for various algorithms with approximation guarantees in CSS. Specifically, if we denote the first $k$ columns of the permutation matrix $\Pi$ as $\Pi_{k}$, then choosing the column subset $C=M\Pi_{k}$ guarantees that $loss(M,C)$ is bounded i.e., $\|M-P_{C}M\|_{F}=\|R_{22}\|_{F}$ (Boutsidis et al., 2009). We tailor Low-RRQR (L-RRQR), originally introduced by Chan and Hansen (1994), to accommodate the two-group fair setting.
@@
Fair pivoting. Note that L-RRQR may introduce unfairness, since we factorize the matrix $M$, without considering the error of the two groups, $A$ and $B$ separately. So the pivoted columns may benefit only one group. To address this, we adapt the pivoting strategy of L-RRQR to benefit the group that suffers the worse reconstruction error. Thus, we perform simultaneous RRQR on $A$ and $B$ and at step-$i$ obtain the corresponding $Q^{A}(i),R^{A}(i)$ and $Q^{B}(i),R^{B}(i)$. We inspect the spectra of both $R^{A}(i)$ and $R^{B}(i)$ and choose the pivot based on the following strategy: we choose the right singular vector $v$ of either $R^{A}(i)$ or $R^{B}(i)$ corresponding to $\max\{\sigma_{1}(R^{A}_{22}(i)),\sigma_{1}(R^{B}_{22}(i))\}$, and select $\Pi$ such that $|(\Pi^{T}v)|_{1}=\|v\|_{\infty}$. The algorithm pseudocode is available in Algorithm 2.
@@ 6 Experiments
Datasets. We use juvenile recidivism data (recidivism) from Catalunya (Tolan et al., 2019) and medical expenditure survey data 2015 (meps(Cohen et al., 2009), as well as various datasets from the UCI-ML repository (Dua and Graff, 2017):“heart-cleveland" (heart), “adult" (adult), “german-credit" (german), “credit-card" (credit), “student performance" (student), “compas-recidivism" (compas), “communities" (communities). Data is processed by removing protected attributes, converting categorical variables to one-hot encoding and normalizing each column to unit $L^{2}$-norm. Group membership is based on Sex, except for communities where group membership is majority white or a non-white community. Dataset statistics are reported in Table 1.
Table 1: Dataset statistics. $m_{A}$ and $m_{B}$ are the number of instances in groups $A$ and $B$, respectively. $n$ is the number of columns. $\gamma$($A$), $\gamma$($B$) is rank of $A$, $B$.
@@ 6.1 Algorithms Evaluation
Figure 1: MinMaxLoss for different values of $c$ and fixed target rank $k$
@@
Table 2: Performance comparison of algorithms.
@@
Table 2 reports the results for the seven largest datasets for various values of $k$. We set $\theta=k-\frac{1}{2}$ in all cases, except meps, where $\theta=\frac{3k}{4}$ to reduce the number of sampled columns. Thus in all datasets (except meps) the resulting number of columns in the first stage, is a $\sqrt{2}$-approximation to the optimal number of columns $c$. Column “$c$" in Table 2 indicates the number of columns returned in the sampling stage. Note that in some cases the number of columns is required to satisfy $\theta$ is significantly large. For each experiment, the best performing algorithm is highlighted in bold.
@@
Figure 2: Comparison of reconstruction error of CSS and FairCSS for groups $A$ and $B$.
@@
Figure 3: Price of fairness.
6.2 Evaluation of the fair CSS objective
We examine the imbalance in reconstruction errors of the “vanilla" CSS objective of Problem 1 and FairCSS-MinMax objective for groups $A$ and $B$. We compute the optimal solution for each objective for various $k$ through exhaustive enumeration. ${\sf opt}(M),{\sf opt}(A)$ and ${\sf opt}(B)$ denote the reconstruction error of vanilla CSS for matrices $M,A$ and $B$, respectively. ${\sf minmax}(M)={\sf minmax}(A,B)$ denote the reconstruction error of matrix $M$. Further, ${\sf fair}(A)$ and ${\sf fair}(B)$ denote the reconstruction errors of FairCSS-MinMax for groups $A$ and $B$, respectively. Lastly, ${\sf fair}(M)$ denote the reconstruction error of $M$ corresponding to the optimal solution of FairCSS-MinMax. We note that brute-force enumeration is expensive, even after parallelization and extensive optimization, so we only report results for heart, student and german datasets.
Fairness of solutions. In Figure 2 we compare ${\sf opt}(A)$, ${\sf opt}(B)$, that is, the optimal solution of vanilla CSS, and the optimal solution of FairCSS-MinMax ${\sf fair}(A),{\sf fair}(B)$. In most instances, we observe that the reconstruction errors of groups are disproportionate in vanilla CSS. However, the degree of imbalance is not uniform across datasets. One source of imbalance could be vastly different group sizes. However, as observed in student, $|A|>|B|$, but ${\sf opt}(A)<{\sf opt}(B)$. This further supports the need for sophisticated approaches to fairness in CSS, beyond mere normalization.
Price of fairness. Next, we verify how the fairness objective influences the quality of the solution in terms of reconstruction error. In Figure 3, we report the optimal solution of vanilla CSS and FairCSS-MinMax to assess the trade-off between fairness and reconstruction error. This analysis quantifies the extent to which we sacrifice reconstruction error to achieve our fairness objective. In most cases, we observe no significant difference in reconstruction error between vanilla CSS and FairCSS-MinMax, that is, ${\sf opt}(M)$ and ${\sf fair}(M)$, even though the value of ${\sf minmax}(M)$ is significantly higher than ${\sf opt}(M)$. Note that, the observations cannot be generalized across all datasets, and we cannot conclusively claim that there is no trade-off between fairness and reconstruction error. For instance, in the case of the heart dataset, we observe a significant difference between values of ${\sf opt}(M)$ and ${\sf fair}(M)$ for $k=5$.
7 Conclusion
@@ Acknowledgements
Suhas Thejaswi acknowledges support from the European Research Council (ERC) under the European Union’s Horizon $2020$ research and innovation program (grant agreement No. $945719$) and the European Unions’s SoBigData++ Transnational Access Scholarship. Antonis Matakos acknowledges support from the Academy of Finland through the grant "Model Management Systems: Machine learning meets Database Systems"- MLDB (32511).
@@ Appendix B Decay of Leverage Scores
Figure 4: Leverage scores of $A$ and $B$ for Table 2